Please help me solve this equation:
(2/3)+(5/8)/(1/2)-(3/8)

Answer by whty89(20)

first you want to get rid of those fractions. i hate fractions
(2/3)+(5/8)/(1/2)-(3/8)
1.) multiply the whole thing by the LCM in this case 24.
that will give you (2/3 x 24=16)+(5/8 x 24=15)/(1/2 x 24=12)-(3/8 x 24=9)
2.)16+(15/12)-9 Order of operations PEMDAS divide first

3.)16+(1.25)-9 ADD and subtract from left to right.
17.25-8
8.25